Two particles are projected from the ground simultaneously with speed 20m/s and 9m/s at angle 30 and 60 with horizontal in the same direction. the maximum distance between them till both of them strike the ground is
 

Grade:11

1 Answers

Vikas TU
7 years ago
Find the range in both cases and then subtract it.
  1. R =u^2sin(2@)/g = 400*sin60/g = 200root3/g
  2. R2 = v^2sin(2$)/g = 81*sin120/g = 81/2root2/g
Distaces between = R1 -R2 = 200root3/g – 81/2root2/g